OpenStudy (anonymous):

Determine the minimum value of the function f(x)=4^x -8. I found the derivative f'(x)-4^x ln4. Would I do f'(x)=0 and solve for x?

OpenStudy (anonymous):

\[4^{x} -8\]

hartnn (hartnn):

yes, f'(x) = 0 gives you extreme points

OpenStudy (anonymous):

Ok I did that and got 0=4^x ln4 and that would turn to 0=4^x. would would i isolate x?

hartnn (hartnn):

can 4^x ever be = 0 ?

hartnn (hartnn):

for any value of x ?

OpenStudy (anonymous):

nope

hartnn (hartnn):

then this function does have any minimum :)

OpenStudy (anonymous):

Oh ok! thanks!

hartnn (hartnn):

or you can say, it has a minimum at \(\Large -\infty \) which is same as saying it does not have a minimum \(4^{-\infty } = 0 \)

hartnn (hartnn):

and welcome ^_^